Calculate the vital capacity when the Tidal Volume is 1.0 liters, the Inspiratory Reserve is 2.0 liters, the Expiratory Reserve is 1.5 liters, and the Residual Volume is 2.5 liters.
Answer #1

Vital Capacity (VC) is the total volume of air that can be moved out from the lungs after maximum inhalation during respiration and formula to calculate VC is-

VC = TV + ERV + IRV

TV- tidal volume

ERV- expiratory reserve volume

IRV- inspiratory reserve volume

There is no role of residual volume in calculating TV. Residual volume is the amount of air that remains in the lungs after respiration and is present in the lungs all the time, without residual volume lungs will get collapsed.

So put all the values given in the question-

TV = 1 L

IRV = 2 L

ERV = 1.5 L

VC = 2+1+1.5 = 4.5 L

So Vital Capacity is 4.5 L.